# L2R: Low-Rank and Lipschitz-Controlled Routing for Mixture-of-Experts

### v1

@@ Abstract
Mixture-of-Experts (MoE) models scale neural networks by conditionally activating a small subset of experts, where the router plays a central role in determining expert specialization and overall model performance. However, many modern MoE systems still adopt linear routers in raw high-dimensional representation spaces, where representation mismatch, angular concentration, and scale-sensitive scoring can jointly undermine routing discriminability and stable expert specialization. In this work, we propose Low-rank & Lipschitz-controlled Routing (L2R), a unified routing framework that reshapes both the routing space and scoring geometry. L2R performs expert assignment in a shared low-rank latent routing space and introduces Saturated Inner-Product Scoring (SIPS) to explicitly control the Lipschitz behavior of routing functions, yielding smoother and more stable routing geometry. In addition, L2R incorporates a parameter-efficient multi-anchor routing mechanism to enhance expert expressiveness. Extensive experiments on a large-scale language MoE model and a vision MoE setting on ImageNet demonstrate that L2R consistently improves routing stability, expert specialization, and overall model performance.
Deep Learning
Email: {yang, togo, guang, ogawa, mhaseyama}@lmd.ist.hokudai.ac.jp
@@ 1 Introduction
Mixture-of-Experts (MoE) has emerged as a leading paradigm for scaling deep neural networks across a wide range of domains, including large language models (LLMs) (Muennighoff et al., 2025; Dai et al., 2024; Jiang et al., 2024; Yang et al., 2024) and vision backbones (Riquelme et al., 2021; Chen et al., 2023; Yang et al., 2025). By enabling conditional computation, MoE decouples model capacity from inference cost and is designed to realize meaningful expert specialization (Cai et al., 2025). Central to achieving this objective is the router. As the mechanism that assigns tokens to experts under sparse and dynamic activation, the router effectively defines the functional topology of the model, governing how capacity is allocated and coordinated. Consequently, the quality of routing is a fundamental factor that determines the expressiveness, stability, and upper performance bound of MoE models (Dikkala et al., 2023).
Figure 1: Comparison between the linear router (left) and the proposed L2R framework (right). (a) Routing space: Tokens are projected from high-dimensional raw representations into a low-rank routing space as $\bm{q}$, where experts are represented by learnable anchors $\bm{k}$. (b) Scoring mode: Compared to dot-product scoring, SIPS reshapes the score landscape into a bounded and smoother geometry, improving routing stability and expert specialization.
Despite its central role, routing in modern MoE models, including OLMoE (Muennighoff et al., 2025), Mixtral (Jiang et al., 2024), Qwen2-MoE (Yang et al., 2024), and DeepSeek-MoE (Dai et al., 2024), still relies on linear routers operating directly on raw token representations. Recent efforts started to revisit this design from different perspectives. Router Upcycling (Ran et al., 2025) proposes a mixture-of-routers paradigm by initializing multiple routers from attention heads, Yuan 2.0-M32 (Wu et al., 2024a) introduces attention-based routing to enrich routing capacity, and X-MoE (Chi et al., 2022) reveals representation collapse and token clustering induced by high-dimensional gating.
@@
Despite these advances, modern MoE routing still exhibits two fundamental limitations: 1) Raw-space routing and high-dimensional concentration: Routing is typically performed in raw, high-dimensional representation spaces that are optimized for representation learning rather than expert discrimination, resulting in a fundamental representation mismatch. Moreover, in high-dimensional spaces, similarities tend to concentrate (Cai et al., 2013), reducing effective separability among experts (Chi et al., 2022). 2) Scale-sensitive dot-product scoring: Standard dot-product logits entangle directional alignment with feature scale, making routing highly sensitive to magnitude variations and encouraging norm-dominated competition among expert anchors, which introduces ambiguity in expert selection and weakens expert specialization.
@@
To address these issues, we propose Low-rank & Lipschitz-controlled Routing (L2R), a unified routing framework with three components: a low-rank routing space for representation decoupling, a Lipschitz-controlled scoring function for stable expert selection, and a parameter-efficient multi-anchor mechanism for enhanced expert expressiveness. First, instead of performing expert selection directly in the raw high-dimensional representation space, tokens are mapped into a shared low-rank routing space that produces routing-specific queries. Experts are represented by learnable anchors in this latent space, and routing is performed by matching queries to anchors. This design decouples routing from representation learning, compresses the search space for expert positioning, and alleviates geometric concentration in high-dimensional routing. Second, Saturated Inner-Product Scoring (SIPS) is introduced to bound the sensitivity of routing scores to representation scale and perturbations, thereby explicitly controlling the Lipschitz behavior of the routing function. By reshaping the score landscape into a smoother and more stable geometry, SIPS reduces routing ambiguity and promotes robust expert specialization. Finally, we develop a parameter-efficient multi-anchor routing mechanism that associates each expert with multiple anchors, allowing a single expert to capture diverse semantic views without incurring prohibitive overhead.
@@
We extensively evaluate the proposed L2R framework across diverse MoE settings to demonstrate its generality and effectiveness. Specifically, we conduct from-scratch training on large-scale language MoE models based on OLMoE (Muennighoff et al., 2025) and evaluate vision MoE architectures on ImageNet using ViT backbones. Across both domains, our method consistently improves routing stability, expert specialization, and overall model performance, validating its effectiveness in both language and vision domains.
@@
  • We identify two fundamental limitations of linear MoE routing, namely routing in raw high-dimensional representation spaces and scale-sensitive dot-product scoring, which lead to routing ambiguity and insufficient expert specialization.
@@
  • We validate L2R across both language and vision MoE settings, including large-scale language models and vision backbones on ImageNet, consistently improving routing stability, expert specialization, and overall performance.
2 Related Work
Router Design in Sparse MoE.
Most modern sparse MoE systems still rely on raw-space linear routing, where expert logits are produced by a single linear projection in the backbone representation space, e.g., in OLMoE (Muennighoff et al., 2025), Mixtral (Jiang et al., 2024), Qwen2-MoE (Yang et al., 2024), and DeepSeek-MoE (Dai et al., 2024), as well as vision MoE variants such as V-MoE (Riquelme et al., 2021), Mod-Squad (Chen et al., 2023), and ASE (Yang et al., 2025). Recent studies revisit router design from multiple angles: improving routing expressiveness (e.g., Router Upcycling (Ran et al., 2025) and attention-based routing in Yuan 2.0-M32 (Wu et al., 2024a)), and improving stability and utilization via auxiliary objectives and routing variants (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022). Despite these advances, the routing score is still typically produced by linear scoring in a high-dimensional space, which leaves the scoring geometry sensitive to scale and high-dimensional concentration.
Geometry-Aware Latent-Space Routing.
Another line of work performs routing in a latent space and adopts angle-aware similarities (Chi et al., 2022; Nguyen et al., 2025; Wu et al., 2024b). X-MoE (Chi et al., 2022) projects representations to a lower-dimensional space and uses cosine scoring to alleviate collapse in high-dimensional gating. Different from prior works, our study emphasizes two structural limitations that jointly weaken discriminability under raw-space routing: representation mismatch and angular concentration. We further observe that common rank heuristics (e.g., scaling rank with the number of experts) can still yield an angle-concentrated routing geometry in practice, motivating an ultra-low yet consistent rank across settings. Within this latent-space routing paradigm, L2R introduces SIPS, a Lipschitz-controlled scoring geometry that retains magnitude cues while explicitly bounding their effective influence (instead of discarding magnitude as cosine scoring), and improves expressiveness via lightweight multi-anchor heads.
@@ 3.1 Preliminaries
MoE models scale network capacity by replicating a sub-module, such as the feed-forward network (FFN) in a Transformer block, into a set of $N$ parallel experts and activating only a sparse subset per input token (Cai et al., 2025). Given an input representation $\bm{x}\in\mathbb{R}^{d}$, a router $g(\cdot)$ first produces routing logits:
@@
where $\mathcal{T}$ denotes the indices of the selected experts. The routing scores jointly determine expert activation and their relative contributions, thereby defining the effective mixture geometry and influencing expert specialization (Dikkala et al., 2023).
@@
In most existing MoE architectures (Muennighoff et al., 2025; Jiang et al., 2024; Yang et al., 2024; Dai et al., 2024), the router $g(\cdot)$ is implemented as a dense linear transformation:
@@ 3.2 Low-Rank Routing
Backbone representations $\bm{x}$ are primarily optimized for representation learning and downstream task prediction, rather than for expert discrimination. Consequently, the router is forced to operate in a passive feature space that is not explicitly shaped for separating tokens into expert-specific clusters. Especially in LLMs, semantic factors in such representations are often highly entangled due to superposition effects (Elhage et al., 2022), and token distributions are not encouraged to organize into well-separated expert-aware regions. This representation mismatch makes it difficult for a linear router to induce stable and semantically coherent expert assignments.
Figure 2: Variance of pairwise cosine similarity in routing spaces. We compare the layer-averaged variance of pairwise token cosine similarities in the latent routing space for Linear, X-MoE (Chi et al., 2022), and L2R-SIPS applied to OLMoE (Muennighoff et al., 2025). For L2R-SIPS, we use rank r=2r{=}2, hence the isotropic reference corresponds to the 2D value ($0.5$). While Linear and X-MoE exhibit near-zero variance, L2R-SIPS attains a much higher variance, indicating substantially improved angular diversity.
@@
Linear routing typically operates in a high-dimensional regime with $d\gg N$ (e.g., $d=4,096$ with $N\leq 64$(Cai et al., 2025). In such spaces, angular similarities tend to concentrate (Cai et al., 2013), causing random or weakly structured vectors to become nearly orthogonal. Figure 2 provides direct empirical evidence of this phenomenon in OLMoE (Muennighoff et al., 2025): backbone token representations for Linear routers exhibit near-zero variance in pairwise cosine similarity, indicating a highly concentrated angular geometry. As a result, routing logits across experts have poorly separated margins, leading to ambiguous expert selection. Moreover, the large dimensionality of the routing space significantly enlarges the search space for expert anchors, making anchor optimization increasingly dependent on initialization and hindering effective expert positioning.
@@
This formulation yields two immediate benefits: First, the shared projection $\bm{W}_{q}$ explicitly learns a routing-oriented subspace, transforming backbone features into a space tailored for expert discrimination and mitigating raw-space representation mismatch. Second, routing in the compressed space $\mathbb{R}^{r}$ substantially reduces geometric concentration (as presented in Figure 2) and the anchor search space, enabling more stable anchor positioning and more discriminative expert specialization.
@@ 3.3 Saturated Inner-Product Scoring (SIPS)
Cosine-style scoring like X-MoE (Chi et al., 2022) stabilizes routing by removing magnitude factors, but it also discards a potentially informative signal for adaptive gating. In contrast, SIPS rebalances scale and angle: it preserves magnitude information for adaptive scoring, while explicitly bounding its influence so that routing remains stable and controllable. Concretely, SIPS factorizes each logit into a bounded magnitude term and an angular term:
@@
(a) Dot Product Scoring
(b) SIPS Scoring
Figure 3: Score landscapes under fixed expert anchor. Heatmaps visualize routing logits as a function of query location $\bm{q}=(Q_{x},Q_{y})$ with a fixed anchor $\bm{k}=[2,2]$. Standard dot-product yields a linear half-space separation with unbounded magnitude effects, while SIPS reshapes the landscape into a bounded, angle-sensitive geometry that is more amenable to stable routing.
@@
We bound query magnitudes through a normalize-then-saturate design. Normalization is applied first to stabilize the distribution of query norms: for vision models, we apply BatchNorm to the scalar norm $\|\bm{q}\|$, while for LLMs we apply RMSNorm (Zhang and Sennrich, 2019) to the router input before the projection $\bm{W}_{q}$. We then define a bounded query magnitude transform:
@@
Here, $\beta\in[0,1]$ controls how much magnitude participates in expert selection ($\beta=0$ removes magnitude effects), and $\gamma$ sets the overall logit scale (equivalently, it rescales logits and modulates the effective softmax temperature). This design mitigates excessively small logits for low-magnitude queries, while $\tanh(\cdot)$ limits extreme magnitudes and prevents outliers from inducing overly sharp routing, hence improving routing stability.
We provide an additional 3D score-landscape comparison across dot-product, cosine, and SIPS with varying $\beta$ in Appendix A, which further illustrates the near-origin smoothness and controlled radial growth induced by query-magnitude saturation.
@@
A key effect of SIPS is to explicitly upper-bound the logit sensitivity to scale variations, yielding a smoother and more controllable scoring geometry. Recall that in Equation (8), $z_{i}=\phi(\|\bm{q}\|)\,\psi(\|\bm{k}_{i}\|)\,\cos\theta_{i}$, the angular term satisfies $\cos\theta_{i}\in[-1,1]$ and is therefore intrinsically bounded. Consequently, scale-induced instability in dot-product scoring primarily arises from the magnitude factors $\phi(\|\bm{q}\|)$ and $\psi(\|\bm{k}_{i}\|)$. SIPS addresses this by (i) bounding the query scale via $\phi(\|\bm{q}\|)$ (Equation (9)) and (ii) compressing anchor norms around $1$ via $\psi(\|\bm{k}_{i}\|)$ (Equation (11)). These designs not only bound the attainable logit magnitude, but also bound the slope of $z_{i}$ with respect to $\|\bm{q}\|$ and $\|\bm{k}_{i}\|$, which directly translates into a smaller Lipschitz constant of the scoring map along the radial directions. Geometrically, this reshapes the fixed-anchor score field from an unbounded linear plane (Figure 3(a)) to a direction-selective landscape with controlled radial growth (Figure 3(b)), mitigating norm-driven ambiguity and improving the stability of expert selection. Formal gradient-sensitivity analysis and the resulting Lipschitz bounds for Equation (8) are provided in Appendix B.

### v2

@@ Abstract
Mixture-of-Experts (MoE) models scale neural networks by conditionally activating a small subset of experts, where the router plays a central role in determining expert specialization and overall model performance. However, many modern MoE systems still adopt linear routers in raw high-dimensional representation spaces, where representation mismatch, angular concentration, and scale-sensitive scoring can jointly undermine routing discriminability and stable expert specialization. In this work, we propose Low-rank & Lipschitz-controlled Routing (L2R), a unified routing framework that reshapes both the routing space and scoring geometry. L2R performs expert assignment in a shared low-rank latent routing space and introduces Saturated Inner-Product Scoring (SIPS) to explicitly control the Lipschitz behavior of routing functions, yielding smoother and more stable routing geometry. In addition, L2R incorporates a parameter-efficient multi-anchor routing mechanism to enhance expert expressiveness. Extensive experiments on an OLMoE-based language MoE model and a vision MoE setting on ImageNet demonstrate that L2R consistently improves routing geometry, expert discrimination, and overall model performance. Code will be released.
@@ 1 Introduction
Mixture-of-Experts (MoE) has emerged as a leading paradigm for scaling deep neural networks across a wide range of domains, including large language models (LLMs) Muennighoff et al. (2025); Dai et al. (2024); Jiang et al. (2024); Yang et al. (2024) and vision backbones Riquelme et al. (2021); Chen et al. (2023); Yang et al. (2025). By enabling conditional computation, MoE decouples model capacity from inference cost and is designed to realize meaningful expert specialization Cai et al. (2025). Central to achieving this objective is the router. As the mechanism that assigns tokens to experts under sparse and dynamic activation, the router effectively defines the functional topology of the model, governing how capacity is allocated and coordinated. Consequently, the quality of routing is a fundamental factor that determines the expressiveness, stability, and upper performance bound of MoE models Dikkala et al. (2023).
@@
Despite these advances, modern MoE routing still exhibits two practical limitations: 1) Raw-space routing and high-dimensional concentration: Routing is typically performed in raw, high-dimensional representation spaces that are optimized for representation learning rather than expert discrimination, resulting in a fundamental representation mismatch. Moreover, in high-dimensional spaces, similarities tend to concentrate Cai et al. (2013), reducing effective separability among experts Chi et al. (2022). 2) Scale-sensitive dot-product scoring: Standard dot-product logits entangle directional alignment with feature scale, making routing highly sensitive to magnitude variations and encouraging norm-dominated competition among expert anchors, which introduces ambiguity in expert selection and weakens expert discrimination.
Figure 1: Comparison between the linear router (left) and the proposed L2R framework (right). (a) Routing space: Tokens are projected from high-dimensional raw representations into a low-rank routing space as $\bm{q}$, where experts are represented by learnable anchors $\bm{k}$. (b) Scoring mode: Compared to dot-product scoring, SIPS reshapes the score landscape into a bounded and smoother geometry, improving routing stability and expert discrimination.
@@
To address these issues, we propose Low-rank & Lipschitz-controlled Routing (L2R), a unified routing framework with three components: a low-rank routing space for representation decoupling, a Lipschitz-controlled scoring function for stable expert selection, and a parameter-efficient multi-anchor mechanism for enhanced expert expressiveness. First, instead of performing expert selection directly in the raw high-dimensional representation space, tokens are mapped into a shared low-rank routing space that produces routing-specific queries. Experts are represented by learnable anchors in this latent space, and routing is performed by matching queries to anchors. This design decouples routing from representation learning, compresses the search space for expert positioning, and alleviates geometric concentration in high-dimensional routing. Second, Saturated Inner-Product Scoring (SIPS) is introduced to bound the sensitivity of routing scores to representation scale and perturbations, thereby explicitly controlling the Lipschitz behavior of the routing function. By reshaping the score landscape into a smoother and more stable geometry, SIPS reduces routing ambiguity and promotes robust expert discrimination. Finally, we develop a parameter-efficient multi-anchor routing mechanism that associates each expert with multiple anchors, allowing a single expert to capture diverse semantic views without incurring prohibitive overhead.
@@
We extensively evaluate the proposed L2R framework across diverse MoE settings to demonstrate its generality and effectiveness. Specifically, we conduct from-scratch training on large-scale language MoE models based on OLMoE Muennighoff et al. (2025) and evaluate vision MoE architectures on ImageNet using ViT backbones. Across both domains, our method consistently improves routing geometry, expert discrimination, and overall model performance, validating its effectiveness in both language and vision domains. We provide an expanded discussion of related work in Appendix A.
@@
  • We identify two practical limitations of linear MoE routing, namely routing in raw high-dimensional representation spaces and scale-sensitive dot-product scoring, which lead to routing ambiguity and insufficient expert specialization.
@@
  • We validate L2R across both language and vision MoE settings, including OLMoE-based language models and vision backbones on ImageNet, consistently improving routing geometry, expert discrimination, and overall performance.
@@ 2.1 Preliminaries
MoE models scale network capacity by replicating a sub-module, such as the feed-forward network (FFN) in a Transformer block, into a set of $N$ parallel experts and activating only a sparse subset per input token Cai et al. (2025). Given an input representation $\bm{x}\in\mathbb{R}^{d}$, a router $g(\cdot)$ first produces routing logits:
@@
where $\mathcal{T}$ denotes the indices of the selected experts. The routing scores jointly determine expert activation and their relative contributions, thereby defining the effective mixture geometry and influencing expert specialization Dikkala et al. (2023).
@@
In most existing MoE architectures Muennighoff et al. (2025); Jiang et al. (2024); Yang et al. (2024); Dai et al. (2024), the router $g(\cdot)$ is implemented as a dense linear transformation:
@@ 2.2 Low-Rank Routing
Backbone representations $\bm{x}$ are primarily optimized for representation learning and downstream task prediction, rather than for expert discrimination. Consequently, the router is forced to operate in a passive feature space that is not explicitly shaped for separating tokens into expert-specific clusters. Especially in LLMs, semantic factors in such representations are often highly entangled due to superposition effects Elhage et al. (2022), and token distributions are not encouraged to organize into well-separated expert-aware regions. This representation mismatch makes it difficult for a linear router to induce stable and semantically coherent expert assignments.
Figure 2: Layer-averaged variance of pairwise token cosine similarities in the routing space.
@@
Linear routing typically operates in a high-dimensional regime with $d\gg N$ (e.g., $d=4,096$ with $N\leq 64$Cai et al. (2025). In such spaces, angular similarities tend to concentrate Cai et al. (2013), causing random or weakly structured vectors to become nearly orthogonal. Figure 2 compares the variance of pairwise token cosine similarities in the latent routing spaces of Linear, X-MoE Chi et al. (2022), and L2R-SIPS on OLMoE Muennighoff et al. (2025). It provides direct empirical evidence of this phenomenon: backbone token representations for Linear routers exhibit near-zero variance in pairwise cosine similarity, indicating a highly concentrated angular geometry. As a result, routing logits across experts have poorly separated margins, leading to ambiguous expert selection. Moreover, the large dimensionality of the routing space significantly enlarges the search space for expert anchors, making anchor optimization increasingly dependent on initialization and hindering effective expert positioning.
@@
This formulation yields two immediate benefits: First, the shared projection $\bm{W}_{q}$ explicitly learns a routing-oriented subspace, transforming backbone features into a space tailored for expert discrimination and mitigating raw-space representation mismatch. Second, routing in the compressed space $\mathbb{R}^{r}$ substantially reduces geometric concentration and the anchor search space, enabling more stable anchor positioning and more discriminative expert selection. Figure 2 empirically supports this effect: with rank r=2r{=}2, L2R-SIPS yields much higher pairwise-cosine variance than Linear and X-MoE, approaching the isotropic 2D reference value of $0.5$ and indicating improved angular diversity.
@@ 2.3 Saturated Inner-Product Scoring (SIPS)
(a) Dot Product Scoring
(b) SIPS Scoring
Figure 3: Score landscapes under fixed expert anchor. Heatmaps visualize routing logits as a function of query location $\bm{q}=(Q_{x},Q_{y})$ with a fixed anchor $\bm{k}=[2,2]$.
@@
Cosine-style scoring like X-MoE Chi et al. (2022) stabilizes routing by removing magnitude factors, but it also discards a potentially informative signal for adaptive gating. In contrast, SIPS rebalances scale and angle: it preserves magnitude information for adaptive scoring, while explicitly bounding its influence so that routing remains stable and controllable. Concretely, SIPS factorizes each logit into a bounded magnitude term and an angular term:
@@
We bound query magnitudes through a normalize-then-saturate design. Normalization is applied first to stabilize the distribution of query norms: for vision models, we apply BatchNorm to the scalar norm $\|\bm{q}\|$, while for LLMs we apply RMSNorm Zhang and Sennrich (2019) to the router input before the projection $\bm{W}_{q}$. We then define a bounded query magnitude transform:
@@
Here, $\beta\in[0,1]$ controls how much magnitude participates in expert selection ($\beta=0$ removes magnitude effects), and $\gamma$ sets the overall logit scale (equivalently, it rescales logits and modulates the effective softmax temperature). This design mitigates excessively small logits for low-magnitude queries, while $\tanh(\cdot)$ limits extreme magnitudes and prevents outliers from inducing overly sharp routing, hence improving routing stability. We provide an additional 3D score-landscape comparison across dot-product, cosine, and SIPS with varying $\beta$ in Appendix D, which further illustrates the near-origin smoothness and controlled radial growth induced by query-magnitude saturation.
@@
A key effect of SIPS is to explicitly upper-bound the logit sensitivity to scale variations, yielding a smoother and more controllable scoring geometry. Recall that in Equation (8), $z_{i}=\phi(\|\bm{q}\|)\,\psi(\|\bm{k}_{i}\|)\,\cos\theta_{i}$, the angular term satisfies $\cos\theta_{i}\in[-1,1]$ and is therefore intrinsically bounded. Consequently, scale-induced instability in dot-product scoring primarily arises from the magnitude factors $\phi(\|\bm{q}\|)$ and $\psi(\|\bm{k}_{i}\|)$. SIPS addresses this by (i) bounding the query scale via $\phi(\|\bm{q}\|)$ (Equation (9)) and (ii) compressing anchor norms around $1$ via $\psi(\|\bm{k}_{i}\|)$ (Equation (11)). These designs not only bound the attainable logit magnitude, but also bound the slope of $z_{i}$ with respect to $\|\bm{q}\|$ and $\|\bm{k}_{i}\|$, which directly translates into a smaller Lipschitz constant of the scoring map along the radial directions. Geometrically, this reshapes the fixed-anchor score field from an unbounded linear plane (Figure 3(a)) to a direction-selective landscape with controlled radial growth (Figure 3(b)), mitigating norm-driven ambiguity and improving the stability of expert selection. Formal gradient-sensitivity analysis and the resulting Lipschitz bounds for Equation (8) are provided in Appendix E.